P14.6A. (LO 2, 3, 4 ) AP The ledger of Zug Limited at October 31, 2021, contains the following summary data:

Correct error from prior period; prepare statement of retained earnings.

Cash dividends—common

$  120,000

Operating expenses

$929,000

Common shares

650,000

Interest expense

54,000

Depreciation expense

87,000

Retained earnings, November 1, 2020

575,000

Service revenue

1,476,000

Your analysis reveals the following additional information:

1. The company has a 25% income tax rate.

2. On March 19, 2021, Zug discovered an error made in the previous fiscal year. A $57,000 payment of a note payable had been recorded as interest expense.

3. On April 10, 2021, common shares costing $75,000 were reacquired for $97,500. This is the first time the company has reacquired common shares.

Instructions

a. Prepare a journal entry to correct the prior period error.

b. Prepare the journal entry to record the reacquisition of common shares.

c. Calculate profit for the year ended October 31, 2021.

d. Prepare the statement of retained earnings for the company for the year ended October 31, 2021.
Answer #1

a. Journal Entry to correct prior period error:

In last year, $ 57000 has been reported as Interest expense, because of which Profit would have been less by $ 57,000 and accordingly "Retained Earnings" account has been reported less by $ 57,000. Hence now we need to increase the Retained earnings account and accordingly the correct account (i.e) Note payable has to be accounted.

Thus, entry will be

Dr. Note Payable account 57,000

Cr. Retained Earnings account 57,000

(Being prior period adjustment entry made to rectify expenses and account for note payable)

b. Journal Entry to record reacquisition of common shares - Par value Method:

Reacquisition of common shares - Company purchase its own shares instead of issuing to any outsiders/beneficiaries.

Lets apply Par value method of accounting for reacquisition of common shares (i.e.) par value of share $75000 But reacquired for $ 97,500 and hence the difference has to be accounted as "Paid-in capital". Treasury stock is the account used for accounting any reacquisition of common share and this will form part of "Equity" schedule in Balance sheet.

Thus entry will be

Dr. Treasury stock 75,000

Dr. Additional Pain-in Capital 22,500

Cr. Cash 97,500

(Being reacquisition of common shares accounted)

c. Profit for year ended October 31, 2021 =$ 304,500 and workings as below:

Income Statement for year ending October 31, 2021
Description Ref Amount ($)
Service Revenue A       1,476,000
Less: Operating Expense B           929,000
Net profit / Earnings before depreciation, interest &Taxes C=A-B           547,000
Less: Depreciation Expense D             87,000
Less: Interest Expense E             54,000
Net Income before Taxes F=C-D-E           406,000
Less: Tax @ 25% G=F*25%           101,500
Net Income after taxes H=F-G           304,500

d. Statement of Retained Earnings for year ended October 31, 2021:

Statement of Retained Earnings:
Description Ref Amount ($)
Opening Balance as on Nov 1, 2020 A           575,000
Add: Current Year Profit B           304,500
Add: Prior Period Adjustment C             57,000
Less: Cash dividends D           120,000
Closing Balance as on Oct 31, 2021 A+B+C-D           816,500